Step 2 in elimination is impossible. Equation is
-4x-2y=-12
4x+8y=-24
Please help

Answer by Alan3354(69443)

-4x-2y=-12
4x+8y=-24
----------- Add
6y = -36
y = -6
-------
-4x-2y=-12
-4x + 12 = -12
x = 6